OpenStudy (saifoo.khan):

@Faye281 : A pyramid has a volume of 20 cubic inches and the area of the base is 15 square inches,What is the height of the pyramid?

OpenStudy (saifoo.khan):

Formula for the volume of pyramid. Vol = 1/3 * area of base* height 20 = 1/3 * 15 * h

mathslover (mathslover):

given volume of pyraimd = 20 cubic inches area of the pyramid = 15 square inches height of the pyramid ?= ? = x inches .. ok ?

OpenStudy (saifoo.khan):

I think i solved the problem already?

mathslover (mathslover):